Stacy has a collection of coins from 5 diffrent countries. Half of the coins are from England, 1/6 are from Mexico, 1/6 are from India, 1/12 are from Italy, and 7 are from Spain. How many coins are in the collection, and how many are from each country?
Answer by richwmiller(17219)

let x= total number of coins
x/2+x/6+x/6+x/12+7=x
x=84
84/2=England
84/6=Mexico and India
84/12=Italy
7 Spain